# Global regularity for solutions of magnetohydrodynamic equations with large initial data

### v6

@@ 1. Introduction
(2.44) $K>3.$
We may represent
(2.45) $\frac{-jf^{\prime}(r)}{\left(r-3j\right)}$ $=$ $\frac{(-Kj+r-r)f^{\prime}(r)}{K\left(r-3j\right)}$
$=$ $\frac{r-M_{q}}{K\left(r-3j\right)}\left[\frac{(r-Kj)f(r)}{\left(r-M_{q}\right)}\right]^{\prime}-\frac{r(r-M_{q})}{K(r-3j)}\left[\frac{f(r)}{(r-M_{q})}\right]^{\prime}$
$-\frac{(r+Kj-M_{q})f(r)}{K(r-3j)(r-M_{q})}.$
We easily verify from (2.42) that
(2.46) $\left[\frac{f(r)}{(r-M_{q})}\right]^{\prime\prime}$ $=$ $\frac{f^{\prime\prime}(r)}{(r-M_{q})}-\frac{2f^{\prime}(r)}{(r-M_{q})^{2}}+\frac{2f(r)}{(r-M_{q})^{3}}$
$\geq$ $-\frac{2}{(r-M_{q})^{2}}\left(f^{\prime}(r)-\frac{f(r)}{(r-M_{q})}\right).$
To estimate the term in parentheses, we invoke (2.33). But we will use a slightly different version of it. Without any loss of generality, we may assume that the second term there is less that or equal to the third term there, i.e.,
$16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}\leq c_{3}\|w\|_{L_{N},Q_{T}}^{s_{2}}\|w\|_{r,Q_{T}}^{\frac{r}{r-M_{q}}}.$
Were this not true, we would have nothing more to prove. Consequently, (2.33) is reduced to
$\|w\|_{\infty,Q_{T}}$ $\leq$ $2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{2}}\|w\|_{r,Q_{T}}^{\frac{r}{r-M_{q}}}.$
Equipped with this, we calculate from (2.41) that
(2.47) $f^{\prime}(r)\leq\ln\left(\|w\|_{\infty,Q_{T}}\right)\leq\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{2}}\right)+\frac{f(r)}{r-M_{q}}.$
Incorporate this into (2.46) to get
(2.48) $\left[\frac{f(r)}{r-M_{q}}\right]^{\prime\prime}$ $\geq$ $\frac{-2\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{2}}\right)}{(r-M_{q})^{2}}.$
We easily infer from (2.38) and (2.39) that there exist two smooth functions $h_{1}(r)$ and $h_{2}(r)$ on $(M_{q},\infty)$ with
$\left[h_{1}(r)+h_{2}(r)\ln\left(\|w\|_{L_{N},Q_{T}}\right)\right]^{\prime\prime}=\frac{2\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{2}}\right)}{(r-M_{q})^{2}}.$
That is,
$h_{1}^{\prime\prime}(r)$ $=$ $\frac{2\ln(2c_{3})}{(r-M_{q})^{2}},$
$h_{2}^{\prime\prime}(r)$ $=$ $\frac{2s_{2}}{(r-M_{q})^{2}}.$
Now (2.48) becomes
(2.49) $\left[\frac{f(r)}{r-M_{q}}+h_{1}(r)+h_{2}(r)\ln\left(\|w\|_{L_{N},Q_{T}}\right)\right]^{\prime\prime}\geq 0\ \ \mbox{on $(M_{q},\infty]$}.$
We further assume
(2.50) $q<r.$
By (2.32), there holds
$q>M_{q}\ \ \mbox{at least for $q$ large}.$
As a result, we may conclude from (2.49) that
$\left[\frac{f(r)}{r-M_{q}}+h_{1}(r)+h_{2}(r)\ln\left(\|w\|_{L_{N},Q_{T}}\right)\right]^{\prime}$
$\geq$ $\frac{\frac{f(r)}{r-M_{q}}+h_{1}(r)+h_{2}(r)\ln\left(\|w\|_{L_{N},Q_{T}}\right)-\frac{f(q)}{q-M_{q}}-h_{1}(q)-h_{2}(q)\ln\left(\|w\|_{L_{N},Q_{T}}\right)}{r-q}.$
For $r>3j$ we may multiply through the inequality by $\frac{r(r-M_{q})}{K(r-3j)}$ to derive
(2.51) $e^{-\frac{r(r-M_{q})}{K(r-3j)}\left[\frac{f(r)}{r-M_{q}}\right]^{\prime}}\leq c\|w\|_{L_{N},Q_{T}}^{\frac{r(r-M_{q})}{K(r-3j)}\left(h_{2}^{\prime}(r)+\frac{h_{2}(q)-h_{2}(r)}{r-q}\right)}\|w\|_{r,Q_{T}}^{-\frac{r^{2}}{K(r-3j)(r-q)}}\|w\|_{q,Q_{T}}^{\frac{r(r-M_{q})q}{K(r-3j)(q-M_{q})(r-q)}}.$
We can handle the first term on the right-hand side of (2.45) in a similar way. Indeed, we compute, with the aid of (2.47), that
(2.52) $\left[\frac{(r-Kj)f(r)}{\left(r-M_{q}\right)}\right]^{\prime\prime}$ $=$ $\frac{(r-Kj)f^{\prime\prime}(r)}{\left(r-M_{q}\right)}+\frac{2(Kj-M_{q})f^{\prime}(r)}{\left(r-M_{q}\right)^{2}}-\frac{2(Kj-M_{q})f(r)}{\left(r-M_{q}\right)^{3}}$
$=$ $\frac{(r-Kj)f^{\prime\prime}(r)}{\left(r-M_{q}\right)}+\frac{2(Kj-M_{q})}{\left(r-M_{q}\right)^{2}}\left[f^{\prime}(r)-\frac{f(r)}{\left(r-M_{q}\right)}\right]$
$\leq$ $\frac{(r-Kj)f^{\prime\prime}(r)}{\left(r-M_{q}\right)}+\frac{2(Kj-M_{q})\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{2}}\right)}{\left(r-M_{q}\right)^{2}}.$
We pick
$r\leq Kj.$
This is possible owing to (2.44). Combing (2.52) with (2.42) yields
$\left[\frac{(r-Kj)f(r)}{\left(r-M_{q}\right)}\right]^{\prime\prime}\leq\frac{2(Kj-M_{q})\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{2}}\right)}{\left(r-M_{q}\right)^{2}}\ \ \mbox{for $r\in(M_{q},Kj]$}.$
As before, there exist two smooth functions $g_{1}(r)$ and $g_{2}(r)$ on $(M_{q},\infty)$ such that
$\left[\frac{(r-Kj)f(r)}{\left(r-M_{q}\right)}-g_{1}(r)-g_{2}(r)\ln\left(\|w\|_{L_{N},Q_{T}}\right)\right]^{\prime\prime}\leq 0\ \ \mbox{on $(M_{q},Kj]$}.$
Remember $q\in(M_{q},r)$. This combined with the preceding inequality implies
$\left[\frac{(r-Kj)f(r)}{\left(r-M_{q}\right)}-g_{1}(r)-g_{2}(r)\ln\left(\|w\|_{L_{N},Q_{T}}\right)\right]^{\prime}$
$\leq$ $\frac{\frac{(r-Kj)f(r)}{\left(r-M_{q}\right)}-g_{1}(r)-g_{2}(r)\ln\left(\|w\|_{L_{N},Q_{T}}\right)-\frac{(q-Kj)f(q)}{\left(q-M_{q}\right)}+g_{1}(q)+g_{2}(q)\ln\left(\|w\|_{L_{N},Q_{T}}\right)}{r-q}.$
For each $r\in(3j,Kj]$ we multiply through the inequality by $\frac{(r-M_{q})}{K(r-3j)}$ to deduce
$e^{\frac{(r-M_{q})}{K(r-3j)}\left[\frac{(r-Kj)f(r)}{\left(r-M_{q}\right)}\right]^{\prime}}\leq c\|w\|_{L_{N},Q_{T}}^{\frac{(r-M_{q})}{K(r-3j)}\left(g_{2}^{\prime}(r)+\frac{g_{2}(q)-g_{2}(r)}{r-q}\right)}\|w\|_{r,Q_{T}}^{\frac{r(r-Kj)}{K(r-3j)(r-q)}}\|w\|_{q,Q_{T}}^{\frac{(r-M_{q})(q-Kj)q}{K(r-3j)(q-M_{q})(r-q)}}.$
Plug (2.45) into (2.43) and then use (2.51) and the preceding estimate to obtain
(2.53) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{3}}\|w\|_{r,Q_{T}}^{G(r)}\|w\|_{q,Q_{T}}^{H(r)},$
@@
$G(r)$ $=$ $\frac{r}{r-3j}-\frac{jr}{(r-3j)(r-q)}-\frac{r(r+Kj-M_{q})}{K(r-3j)(r-M_{q})},$
$H(r)$ $=$ $\frac{b}{\alpha(r-3j)}+\frac{(r-M_{q})(r+q-Kj)q}{K(r-3j)(q-M_{q})(r-q)},$
$s_{3}$ $=$ $s_{1}+\frac{r(r-M_{q})}{K(r-3j)}\left(h_{2}^{\prime}(r)+\frac{h_{2}(q)-h_{2}(r)}{r-q}\right)$
$+\frac{(r-M_{q})}{K(r-3j)}\left(g_{2}^{\prime}(r)+\frac{g_{2}(q)-g_{2}(r)}{r-q}\right).$
We take
$j>\frac{q}{3}\ \ \mbox{and}\ \ \ r=Kj.$
Subsequently, (2.22) and (2.50) are both met. Moreover,
(2.54) $G(Kj)$ $=$ $\frac{K}{K-3}-\frac{3Kj-M_{q}}{(K-3)(Kj-M_{q})}$
$=$ $1-\frac{2M_{q}}{(K-3)(Kj-M_{q})},$
$H(Kj)$ $=$ $\frac{b}{\alpha(K-3)j}+\frac{(Kj-M_{q})q^{2}}{K(K-3)(q-M_{q})j(Kj-q)}$
$=$ $\frac{1}{(K-3)j}\left[\frac{q-2L_{N}-2\alpha qj}{\alpha(q-L_{N})}+\frac{(Kj-M_{q})q^{2}}{K(q-M_{q})(Kj-q)}\right]$
$=$ $\frac{1}{(K-3)j}\left[\frac{q-2L_{N}-2\alpha qj}{\alpha(q-L_{N})}+\frac{q^{2}}{K(q-M_{q})}+\frac{q^{2}}{K(Kj-q)}\right].$
Here we have used (2.21). The term between the two parentheses is a decreasing function of $j$ on $\left(\frac{q}{K},\infty\right)$. We evaluate it at $j=\frac{q}{3}$ to deduce
$\left.\frac{q-2L_{N}-2\alpha qj}{\alpha(q-L_{N})}+\frac{q^{2}}{K(q-M_{q})}+\frac{q^{2}}{K(Kj-q)}\right|_{j=\frac{q}{3}}$
$=$ $\frac{3q-6L_{N}-2\alpha q^{2}}{3\alpha(q-L_{N})}+\frac{q^{2}}{K(q-M_{q})}+\frac{3q}{K(K-3)}$
$=$ $-\frac{(2K-9)q}{3(K-3)}+\frac{q-2L_{N}}{\alpha(q-L_{N})}-\frac{2L_{N}q}{3(q-L_{N})}+\frac{M_{q}q}{K(q-M_{q})}$
$<$ $0\ \ \mbox{for $q$ large and $K>\frac{9}{2}$,}$
under which we have
$H(Kj)<0.$
Remember that (2.33) also holds for $r=q$. As a result, we may assume
$\|w\|_{q,Q_{T}}\geq 1.$
Hence,
$\|w\|_{q,Q_{T}}^{H(Kj)}\leq 1.$
Collect this and (2.54) in (2.53) to get
$\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{3}}\|w\|_{r,Q_{T}}^{1-\frac{2M_{q}}{(K-3)(Kj-M_{q})}}.$
Set
$\theta=1-\frac{2M_{q}}{(K-3)(Kj-M_{q})}<1.$
We can choose $K$ so large that
$\theta>0.$
Then (1.5) follows from a suitable application of Young’s inequality (1.10). The proof of Theorem 1.1 is complete. ∎

### v7

@@ 1. Introduction
(2.60) $d\in(M_{q},q).$
Consider the line through $(d,f(d))$ and $(\xi,f(\xi))$. The convexity of $f(s)$ implies
$f(s)$ $\leq$ $f(\xi)+\frac{f(\xi)-f(d)}{\xi-d}(s-\xi)\ \ \mbox{for $s\in(q,\xi)$},$
$f(s)$ $\geq$ $f(\xi)+\frac{f(\xi)-f(d)}{\xi-d}(s-\xi)\ \ \mbox{for $s\in(\xi,r)$.}$
Consequently,
(2.61) $\int_{q}^{\xi}f(s)ds-A\int_{\xi}^{r}f(s)ds$
$\leq$ $f(\xi)(\xi-q)-\frac{(f(\xi)-f(d))(\xi-q)^{2}}{2(\xi-d)}-Af(\xi)(r-\xi)-\frac{A(f(\xi)-f(d))(r-\xi)^{2}}{2(\xi-d)}$
$=$ $f(\xi)\left[\xi-q-A(r-\xi)-\frac{(\xi-q)^{2}+A(r-\xi)^{2}}{2(\xi-d)}\right]+\frac{f(d)\left[(\xi-q)^{2}+A(r-\xi)^{2}\right]}{2(\xi-d)}.$
We pick $A$ so that
(2.62) $\xi-q-A(r-\xi)-\frac{(\xi-q)^{2}+A(r-\xi)^{2}}{2(\xi-d)}=0.$
This yields
(2.63) $A=\frac{\xi-q-\frac{(\xi-q)^{2}}{2(\xi-d)}}{r-\xi+\frac{(r-\xi)^{2}}{2(\xi-d)}}=\frac{(\xi-q)(\xi+q-2d)}{(r-\xi)(r+\xi-2d)}.$
We easily see from (2.60) that $A$ so chosen satisfies (2.57), which, in turn, results in (2.55). We can also confirm this by directly substituting (2.63) into the expression for $h(r)$ in (2.56), thereby obtaining
(2.64) $h(r)$ $=$ $\frac{(\xi-q)\left[-(\xi+q-2d)(r+\xi-2M_{q})+(r+\xi-2d)(\xi+q-2M_{q})\right]}{2(r-M_{q})(r+\xi-2d)}$
$=$ $\frac{(\xi-q)[(d-M_{q})r+M_{q}q-d^{2}]}{(r-M_{q})(r+\xi-2d)}>0.$
The last step here is due to (2.60). Now we incorporate (2.62) and (2.63) into (2.61) to deduce
$\int_{q}^{\xi}f(s)ds-A\int_{\xi}^{r}f(s)ds$ $\leq$ $f(d)\left[\xi-q-A(r-\xi)\right]$
$=$ $\frac{f(d)(\xi-q)(r-q)}{r+\xi-2d}.$
Collecting this in (2.59), we arrive at
$\int_{q}^{r}h(s)f^{\prime}(s)ds+\int_{q}^{r}h^{\prime}(s)f(s)ds$
$\leq$ $\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)\int_{q}^{r}h(s)ds+\frac{f(d)(\xi-q)(r-q)}{r+\xi-2d}.$
Use this in (2.58) to obtain
(2.65) $\frac{f(r)}{r-2j}$ $\leq$ $\frac{1}{h(r)(r-2j)}\left[\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)\int_{q}^{r}h(s)ds+\frac{f(d)(\xi-q)(r-q)}{r+\xi-2d}\right].$
On the other hand, we can infer from (2.46) that
$-\frac{jf^{\prime}(r)}{r-2j}\leq-\frac{jf^{\prime}(q)}{r-2j}\leq-\frac{j(f(q)-f(L_{N}))}{(r-2j)(q-L_{N})}.$
Collect this and (2.65) in (2.48) obtain
(2.66) $\|w\|_{\infty,Q_{T}}\leq 16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{3}}\|w\|_{q,Q_{T}}^{H(r)}\|w\|_{d,Q_{T}}^{\frac{d(\xi-q)(r-q)}{h(r)(r-2j)(r+\xi-2d)}},$
@@
(2.67) $H(r)$ $=$ $\frac{q(M_{q}-j)}{(q-L_{N})(r-2j)}-\frac{jq}{(q-L_{N})\left(r-2j\right)},$
$s_{3}$ $=$ $s_{1}+\frac{jL_{N}}{(q-L_{N})\left(r-2j\right)}+\frac{s_{0}\int_{q}^{r}h(s)ds}{h(r)(r-2j)}.$
In view of (2.60), we can form the interpolation inequality
$\|w\|_{d,Q_{T}}\leq\|w\|_{q,Q_{T}}^{\frac{q(d-L_{N})}{d(q-L_{N})}}\|w\|_{L_{N},Q_{T}}^{\frac{L_{N}(q-d)}{d(q-L_{N})}}.$
Incorporate this into (2.66) to obtain
(2.68) $\|w\|_{\infty,Q_{T}}\leq 16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{3}+\frac{(\xi-q)(r-q)L_{N}(q-d)}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}}\|w\|_{q,Q_{T}}^{H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}}.$
We must verify that
(2.69) $H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}<\frac{q}{q-L_{N}}.$
Plug (2.67) and (2.64) into this to derive
$M_{q}+\frac{(r-q)(d-L_{N})(r-M_{q})}{(d-M_{q})r+M_{q}q-d^{2}}<r.$
In view of (2.43) and (2.60), we can solve this inequality for $r$ to obtain
(2.70) $r<\frac{(d+M_{q}-L_{N})q-d^{2}}{M_{q}-L_{N}}.$
This condition must be consistent with (2.50). That is, there must hold
(2.71) $\frac{(d+M_{q}-L_{N})q-d^{2}}{M_{q}-L_{N}}>q.$
This inequality is an easy consequence of (2.60).
In summary, the order in which we pick our parameters is as follows: We first fix $q$ so that both (2.16) and (2.49) hold. Then pick $d$ as in (2.60), from which (2.71) follows. As a result, we can take
$r\in\left(q,\frac{(d+M_{q}-L_{N})q-d^{2}}{M_{q}-L_{N}}\right).$
This implies both (2.70) and (2.50). Now we can define $A$ as in (2.63) for $\xi\in(q,r)$. Finally, $j$ can be any number in the interval $\left(0,\frac{q}{2}\right)$. Hence, (2.22) is valid.
Without any loss of generality, we may assume
(2.72) $\|w\|_{q,Q_{T}}>1\ \ \mbox{for each $q>M_{q}$.}$
Otherwise, (2.34) would be enough to imply our theorem. Under (2.72), we may suppose
(2.73) $H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}>0.$
Were this not true, (2.72) combined with (2.68) would yield our theorem. We can form the interpolation inequality
$\|w\|_{q,Q_{T}}\leq\|w\|_{\infty,Q_{T}}^{\frac{q-L_{N}}{q}}\|w\|_{L_{N},Q_{T}}^{\frac{L_{N}}{q}}.$
Collect this in (2.68) and keep (2.73) in mind to get
$\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{4}}\|w\|_{\infty,Q_{T}}^{\theta},$
where
$\theta$ $=$ $\frac{(q-L_{N})}{q}\left[H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}\right],$
$s_{4}$ $=$ $s_{3}+\frac{(\xi-q)(r-q)L_{N}(q-d)}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}$
$+\frac{L_{N}}{q}\left[H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}\right].$
According to (2.69) and (2.73), we have
$\theta\in(0,1).$
Then (1.5) follows from a suitable application of Young’s inequality (1.10). The proof of Theorem 1.1 is completed. ∎